# SMReferences ℓ

### v1

@@ Introduction
Quantum integrable models have long provided a unique window into the non‐perturbative behavior of quantum many‐body systems, from the exact solution of one‐dimensional quantum magnets [1] to the emergence of generalized hydrodynamics in far‐from‐equilibrium dynamics [2, 3]. Although a universally accepted definition of quantum integrability is lacking [4, 5], its essential feature is often regarded as the existence of an infinite hierarchy of local conserved quantities, which we adopt as the working definition in this Letter. Such conserved quantities encode much information about the initial state and constrain the dynamics in a way that generically prevents thermalization [6], enabling long-lived memory [7] and anomalous transport phenomena [8]. In contrast, generic quantum many-body systems admit only a few trivial conserved quantities, such as total energy and particle number. Therefore, integrability is a highly non-generic and fine-tuned property, and identifying integrable systems has been one of the central problems in rigorous statistical mechanics.
@@
Traditionally, one establishes integrability by solving the system explicitly; for example, solving the Yang–Baxter equation and constructing a commuting family of local charges through the logarithmic derivative of the transfer matrix [9]. As an alternative, Reshetikhin proposed a simpler criterion [10]: the existence of a single third-order conservation law with a special form. Unlike approaches that rely on additional algebraic structures such as the Yang–Baxter equation or the self-duality [11], this criterion is algorithmically verifiable directly from the Hamiltonian. Since this Reshetikhin condition is a necessary condition for the $R$-matrix with difference form to solve the Yang–Baxter equation, it has long served as a guide in the search for new solvable systems [12, 13, 14, 15, 16, 17, 18]. Remarkably, in all known cases, systems satisfying the Reshetikhin condition have indeed turned out to be integrable. For this reason, it has been widely regarded as a heuristic test of integrability, as proposed by Grabowski and Mathieu [19].
@@
Despite its wide use, however, it has remained an open question whether the Reshetikhin condition alone suffices to ensure integrability, namely the existence of infinitely many commuting local conserved quantities. It has often been argued that higher-order conservation laws and their mutual commutativity require additional algebraic input not captured by the third-order condition alone [14]. A recent preprint [20] claimed to resolve this question; however, its argument contains a critical flaw: a key equation employed in the proof is incorrect (see Supplemental Material [21] for details). In this Letter, we present the first complete and correct proof that any one‑dimensional quantum system satisfying the Reshetikhin condition necessarily gives rise to an infinite sequence of commuting local charges. Specifically, we show that the so-called boost operator [22, 23, 24] allows the iterative generation of higher‑order conservation laws. We further discuss a rigorous foundation for the empirical dichotomy of integrability: either a system possesses infinitely many local conserved quantities, or a few simple at most two-body conserved charges.
Setup and main result
@@
We consider a quantum spin system on an infinite open chain, with sites labeled by $\mathbb{Z}$. Each site $i\in\mathbb{Z}$ has a local Hilbert space $\mathcal{H}_{i}\cong\mathbb{C}^{d}$, where $d<\infty$. We denote by $\mathcal{B}_{i}$ the set of bounded linear operators on $\mathcal{H}_{i}$, and by $\mathcal{B}_{0i}$ the subset of traceless operators. We call an operator at most $n$-local if it can be expressed as a sum of terms acting nontrivially on at most $n$ contiguous sites, and $n$-local if it is not (n1)(n-1)-local.
@@
with $\hat{h}^{(2)}_{i}\in\mathcal{B}_{0i}\otimes\mathcal{B}_{0,i+1}$ and $\hat{h}^{(1)}_{i}\in\mathcal{B}_{0i}$. We assume Hermiticity and translational invariance. The mathematically rigorous definitions of the formal sum in Eq. (1) and of the commutators below are given in Supplemental Material [21].
A key ingredient in our analysis is the boost operator [22, 23, 24].
Definition 1.
Fix an arbitrary translationally invariant one-local operator $\hat{C}=\sum_{i\in\mathbb{Z}}\hat{c}_{i}$. The boost operator associated with the Hamiltonian $\hat{H}$ is defined by
$\hat{B}_{\hat{H}}\coloneqq\sum_{j\in\mathbb{Z}}j(\hat{h}^{(2)}_{j}+\hat{h}^{(1)}_{j})+\hat{C}$. (2)
We set $\hat{Q}^{(2)}\coloneqq\hat{H}$ and recursively $\hat{Q}^{(n+1)}\coloneqq[\hat{B}_{\hat{H}},\hat{Q}^{(n)}]$. By construction, $\hat{Q}^{(n)}$ is at most $n$-local. Although the full structure of $\hat{Q}^{(n)}$ is complicated, its $n$-local component admits the compact expression
@@
For any translationally invariant operator $\hat{Q}$ commuting with $\hat{H}$, the commutator $\hat{B}_{\hat{H}},\hat{Q}$] is again translationally invariant [26, 27]. Indeed, denoting by $\mathcal{T}:\mathcal{B}_{i}\to\mathcal{B}_{i-1}$ the translation operator, one finds $\mathcal{T}(\hat{B}_{\hat{H}})=\hat{B}_{\hat{H}}+\hat{H}$, and hence
@@
$=-\sum_{i\in\mathbb{Z}}[\hat{h}^{(2)}_{i},\hat{h}^{(2)}_{i+1}+\hat{h}^{(1)}_{i+1}+\hat{c}^{(1)}_{i}+\hat{c}^{(1)}_{i+1}$]
$\hskip 28.45274pt-\sum_{i\in\mathbb{Z}}[\hat{h}^{(1)}_{i},\hat{c}^{(1)}_{i}$]
$=-\sum_{i\in\mathbb{Z}}[\hat{h}_{\hat{C},i},\hat{h}_{\hat{C},i+1}],$ (5)
where $\hat{h}_{\hat{C},i}\coloneqq\hat{h}^{(2)}_{i}+\hat{h}^{(1)}_{i}+\hat{c}^{(1)}_{i}-\hat{c}^{(1)}_{i+1}\in\mathcal{B}_{i}\otimes\mathcal{B}_{i+1}$.
@@
Since commutativity of translationally invariant local operators is a local property (see Theorem S10 and Corollary S13 in Supplemental Material [21]), Theorem 2 extends to sufficiently large finite systems. Concretely, for a translationally invariant local operator $\hat{Q}=\sum_{i\in\mathbb{Z}}\hat{q}_{i}$ in the infinite system, we can define its finite-size counterpart $\hat{Q}_{[N]}=\sum_{i=1}^{N}\hat{q}_{i}$ on a chain of length $N$ with the periodic boundary condition. Then we have the following corollary.
@@
Proof of Theorem 2
@@
Let $M\geq 1$. If $\hat{H},\hat{Q}^{(2m+1)}]=0$ holds for all $1\leq m\leq M$, then $\hat{Q}^{(k+1)},\hat{Q}^{(l+1)}]=0$ holds for $k+l\leq 2M+2$.
@@
We proceed by induction on $M$. For $M=1$, the only nontrivial case is $\hat{H},\hat{Q}^{(4)}]=0$. This can be shown from the Jacobi identity as follows:
$[\hat{H},\hat{Q}^{(4)}]$ $=[\hat{H},[\hat{B}_{\hat{H}},\hat{Q}^{(3)}]]$
$=[\hat{Q}^{(3)},[\hat{B}_{\hat{H}},\hat{H}]]+[\hat{B}_{\hat{H}},[\hat{H},\hat{Q}^{(3)}]]$
$=[\hat{Q}^{(3)},\hat{Q}^{(3)}]=0$ (6)
Assume that the claim holds up to $M=M^{\prime}-1\geq 1$, and consider $M=M^{\prime}$. The only nontrivial part to prove is that $\hat{Q}^{(k+1)},\hat{Q}^{(l+1)}]=0$ for $2M^{\prime}+1\leq k+l\leq 2M^{\prime}+2$. First, consider the case $k+l=2M^{\prime}+1$. By the Jacobi identity, all such commutators are proportional to each other:
$[\hat{Q}^{(2)},\hat{Q}^{(2M^{\prime}+1)}]$ $=-[\hat{Q}^{(3)},\hat{Q}^{(2M^{\prime})}]+[\hat{B}_{\hat{H}},[\hat{Q}^{(2)},\hat{Q}^{(2M^{\prime})}]]$
$=-[\hat{Q}^{(3)},\hat{Q}^{(2M^{\prime})}]$
$=[\hat{Q}^{(4)},\hat{Q}^{(2M^{\prime}-1)}]-[\hat{B}_{\hat{H}},[\hat{Q}^{(3)},\hat{Q}^{(2M^{\prime}-1)}]]$
$=[\hat{Q}^{(4)},\hat{Q}^{(2M^{\prime}-1)}]$
@@
$=(-1)^{M^{\prime}-1}[\hat{Q}^{(M^{\prime}+1)},\hat{Q}^{(M^{\prime}+2)}]$ (7)
where we used the induction hypothesis $\hat{Q}^{(k+1)},\hat{Q}^{(2M^{\prime}-k+1)}]=0$. In particular, for $k=1$ the hypothesis already gives $\hat{Q}^{(2)},\hat{Q}^{(2M^{\prime}+1)}]=0$. Hence, $\hat{Q}^{(k+1)},\hat{Q}^{(l+1)}]=0$ holds for all $k+l=2M^{\prime}+1$.
For $k+l=2M^{\prime}+2$, the case $k=1$ follows similarly as in Eq. (6). Moreover, proportionality as in Eq. (7) again holds, so $\hat{Q}^{(k+1)},\hat{Q}^{(l+1)}]=0$ is established for all pairs $(k,l)$ with $k+l=2M^{\prime}+2$. ∎
From this lemma, it suffices to prove Theorem 2 by showing: If $\hat{H},\hat{Q}^{(n)}]=0$ holds for some $n\geq 3$, then $\hat{H},\hat{Q}^{(n+1)}]=0$ also holds. The key step is the following observation.
@@
Consider a finite system of size $N$ with the periodic boundary condition. Note that the commutativity of translationally invariant local operators is a local property, and hence independent of $N$. Thus, for sufficiently large $N$ compared with $n$, the assumption can be replaced with the finite-size relation $\hat{H}_{[N]},[\hat{H}_{[N]},\hat{A}^{(n)}_{[N]}]]=0$ (See Corollary S13 of Supplemental Material [21] for a formal proof). By considering the Hilbert-Schmidt norm of $\hat{H}_{[N]},\hat{A}^{(n)}_{[N]}$], we obtain
@@
Together, these lemmas establish Theorem 2. It should be noted that we cannot dispense with the assumption of translational invariance in Lemma 6, since $\hat{H},[\hat{H},\hat{B}_{\hat{H}}]]=0$ does not necessarily imply that $\hat{Q}^{(3)}=-[\hat{H},\hat{B}_{\hat{H}}]=0$.
Dichotomy on integrability

### v3

# Integrability from a Single Conservation Law in Quantum Spin Chains

@@ I Introduction
Quantum integrable models have long provided a unique window into the non-perturbative behavior of quantum many-body systems, from the exact solution of one-dimensional quantum magnets [1] to the emergence of generalized hydrodynamics in far-from-equilibrium dynamics [2, 3]. Although a universally accepted definition of quantum integrability is lacking [4, 5], its essential feature is often regarded as the existence of an infinite hierarchy of local conserved quantities, which we adopt as the working definition in this paper. Such conserved quantities encode much information about the initial state and constrain the dynamics in a way that generically prevents thermalization [6], enabling long-lived memory [7] and anomalous transport phenomena [8]. In contrast, generic quantum many-body systems admit only a few trivial conserved quantities, such as total energy and particle number. Therefore, integrability is a highly non-generic and fine-tuned property, and identifying integrable systems has been one of the central problems in rigorous statistical mechanics.
@@
Traditionally, one establishes integrability by solving the system explicitly; for example, solving the Yang–Baxter equation and constructing a commuting family of local charges through the logarithmic derivative of the transfer matrix [9]. As an alternative, Reshetikhin proposed a simpler criterion [10]: the existence of a single third-order conservation law with a special form. The equivalent condition had earlier appeared in the work of Jimbo and Miwa [11], where it was formulated as the tangential star-triangle relation. Unlike approaches that rely on additional algebraic structures such as the Yang–Baxter equation or the self-duality [12], this criterion is algorithmically verifiable directly from the Hamiltonian. Since this Reshetikhin condition is a necessary condition for the $R$-matrix with difference form to solve the Yang–Baxter equation, it has long served as a guide in the search for new solvable systems [13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24]. Remarkably, in all known cases, systems satisfying the Reshetikhin condition have indeed turned out to be integrable. For this reason, it has been widely regarded as a heuristic test of integrability, as proposed by Grabowski and Mathieu [25].
@@
Despite its wide use, however, it has remained an open question whether the Reshetikhin condition alone suffices to ensure integrability, namely the existence of infinitely many commuting local conserved quantities. It has often been argued that higher-order conservation laws and their mutual commutativity require additional algebraic input not captured by the third-order condition alone [17]. In this paper, we present the first complete proof that any one‑dimensional quantum system satisfying the Reshetikhin condition necessarily gives rise to an infinite sequence of commuting local charges. Specifically, we show that the so-called boost operator [26, 27, 28] allows the iterative generation of higher‑order conservation laws. We further discuss a rigorous foundation for the empirical dichotomy of integrability: either a system possesses infinitely many local conserved quantities, or a few simple at most two-body conserved charges.
The paper is organized as follows. Section II introduces the precise setup and states our main theorem, which asserts that the Reshetikhin condition implies quantum integrability. In Section III, we provide a rigorous definition of operators on the infinite chain. The technical core of the proof, namely the correspondence between infinite and finite systems, is established in Appendix A. Section IV presents a detailed proof of the main result. Section V discusses the empirical dichotomy of integrability in light of the theorem. Finally, in Section VI, we establish extensions of our result to systems with unconventional boost operators and parameter-dependent Hamiltonians.
II Setup and main result
@@
We consider a quantum spin system on an infinite open chain, with sites labeled by $\mathbb{Z}$. Each site $i\in\mathbb{Z}$ has a local Hilbert space $\mathcal{H}_{i}\cong\mathbb{C}^{d}$, where $d<\infty$. We denote by $\mathcal{B}_{i}$ the set of bounded linear operators on $\mathcal{H}_{i}$, and by $\mathcal{B}_{0i}$ the subset of traceless operators. We call an operator at most $n$-local if it can be expressed as a sum of terms acting nontrivially on at most $n$ contiguous sites, and $n$-local if it is not at most (n1)(n-1)-local.
@@
with $\hat{h}^{(2)}_{i}\in\mathcal{B}_{0i}\otimes\mathcal{B}_{0,i+1}$ and $\hat{h}^{(1)}_{i}\in\mathcal{B}_{0i}$. We assume Hermiticity and translational invariance. The mathematically rigorous definitions of the formal sum in Eq. (1) and of the commutators below are given in Section III.
A key ingredient in our analysis is the boost operator [26, 27, 28]. Fix an arbitrary translationally invariant one-local operator $\hat{C}=\sum_{i\in\mathbb{Z}}\hat{c}_{i}$. The boost operator associated with the Hamiltonian $\hat{H}$ is defined by
$\hat{B}_{\hat{H}}$ $\coloneqq\sum_{j\in\mathbb{Z}}j(\hat{h}^{(2)}_{j}+\hat{h}^{(1)}_{j})+\hat{C}$
$=\sum_{j\in\mathbb{Z}}j\hat{h}_{\hat{C},j},$ (2)
where $\hat{h}_{\hat{C},j}\coloneqq\hat{h}^{(2)}_{j}+\hat{h}^{(1)}_{j}+\hat{c}^{(1)}_{j}-\hat{c}^{(1)}_{j+1}\in\mathcal{B}_{i}\otimes\mathcal{B}_{i+1}$. We set $\hat{Q}^{(2)}\coloneqq\hat{H}$ and recursively $\hat{Q}^{(n+1)}\coloneqq[\hat{B}_{\hat{H}},\hat{Q}^{(n)}]$. By construction, $\hat{Q}^{(n)}$ is at most $n$-local. Although the full structure of $\hat{Q}^{(n)}$ is complicated, its $n$-local component admits the compact expression
@@
For any translationally invariant operator $\hat{Q}$ commuting with $\hat{H}$, the commutator $\hat{B}_{\hat{H}},\hat{Q}$] is again translationally invariant [30, 31]. Indeed, denoting by $\mathcal{T}:\mathcal{B}_{i}\to\mathcal{B}_{i-1}$ the translation map, one finds $\mathcal{T}(\hat{B}_{\hat{H}})=\hat{B}_{\hat{H}}+\hat{H}$, and hence
@@
$=-\sum_{i\in\mathbb{Z}}[\hat{h}_{\hat{C},i},\hat{h}_{\hat{C},i+1}].$ (5)
@@
Since commutativity of translationally invariant local operators is a local property (see Theorem 16 and Corollary 19 in Appendix A), Theorem 1 extends to sufficiently large finite systems. Concretely, for a translationally invariant local operator $\hat{Q}=\sum_{i\in\mathbb{Z}}\hat{q}_{i}$ in the infinite system, we can define its finite-size counterpart $\hat{Q}_{[N]}=\sum_{i=1}^{N}\hat{q}_{i}$ on a chain of length $N$ with the periodic boundary condition. Then we have the following corollary.
@@
III Mathematical Framework
In this section, we introduce rigorous definitions of operators and their commutators on the infinite chain, thereby providing a mathematical justification for the formal manipulations used throughout this paper.
III.1 Operators with finite support
The set of all finite subsets of $\mathbb{Z}$ is denoted by
$\mathcal{P}_{\mathrm{fin}}(\mathbb{Z})\coloneqq\{A\subset\mathbb{Z}\mid|A|<\infty\}.$ (6)
For $A\in\mathcal{P}_{\mathrm{fin}}(\mathbb{Z})$, the set of operators supported on $A$ is defined as $\mathcal{B}_{0}^{A}\coloneqq\otimes_{i\in A}\mathcal{B}_{0i}$. We set $\mathcal{B}_{0}^{\emptyset}\coloneqq\mathbb{C}$. The space of (formal) operators on $\mathbb{Z}$ is defined as
$\mathcal{A}\coloneqq\{\Phi:\mathcal{P}_{\mathrm{fin}}(\mathbb{Z})\ni A\mapsto\Phi(A)\in\mathcal{B}_{0}^{A}\}.$ (7)
We formally write $\Phi=\sum_{A\in\mathcal{P}_{\mathrm{fin}}(\mathbb{Z})}\Phi(A)$ and define its adjoint as $\Phi^{\dagger}\coloneqq\sum_{A\in\mathcal{P}_{\mathrm{fin}}(\mathbb{Z})}\Phi(A)^{\dagger}$.
Next, we define the notion of locality for elements of $\mathcal{A}$. For $A\in\mathcal{P}_{\mathrm{fin}}(\mathbb{Z})\setminus\{\emptyset\}$, the range of $A$ is defined as
$r(A)\coloneqq\max A-\min A+1\in\mathbb{Z}_{\geq 1},$ (8)
and we set $r(\emptyset)\coloneqq 0$. For $\Phi\in\mathcal{A}\setminus\{0_{\mathcal{A}}\}$, its range is defined by
$r(\Phi)\coloneqq\sup\{r(A)\mid\Phi(A)\neq 0_{A}\}\in\mathbb{Z}_{\geq 0}\cup\{\infty\},$ (9)
with $r(0_{\mathcal{A}})\coloneqq 0$. We say that $\Phi$ is $r$-local if $r(\Phi)=r$. If $\Phi(\{i,i+1\})=\hat{h}_{i}^{(2)}$ and $\Phi(\{i\})=\hat{h}_{i}^{(1)}$, then $\Phi$ corresponds to the Hamiltonian with nearest-neighbor interactions and on-site potentials discussed in Section II. This is $2$-local whenever $\hat{h}_{i}^{(2)}\neq 0$.
For $A\in\mathcal{P}_{\mathrm{fin}}(\mathbb{Z})$ and $x\in\mathbb{Z}$, we denote by
$\tau_{A}(x):\mathcal{H}^{\otimes A}\mapsto\mathcal{H}^{\otimes(A+x)}$ (10)
the unitary translation map, where $A+x\coloneqq\{a+x\mid a\in A\}\in\mathcal{P}_{\mathrm{fin}}(\mathbb{Z})$. The translation superoperator $\mathcal{T}(x)$ is defined by
$\mathcal{T}(x)$ $:\mathcal{A}\ni\Phi=\sum_{A\in\mathcal{P}_{\mathrm{fin}}(\mathbb{Z})}\Phi(A)$
$\mapsto\sum_{A\in\mathcal{P}_{\mathrm{fin}}(\mathbb{Z})}\tau_{A}(x)^{\dagger}\Phi(A+x)\tau_{A}(x)\in\mathcal{A}.$ (11)
We also write $\mathcal{T}=\mathcal{T}(1)$. This coincides with the translation map introduced in Section II.
We now introduce some subsets of $\mathcal{A}$:
$\mathcal{A}^{\mathrm{sa}}$ $\coloneqq\{\Phi\in\mathcal{A}\mid\Phi^{\dagger}=\Phi\},$ (12)
[…]
$\mathcal{A}^{\mathrm{inv}}$ $\coloneqq\{\Phi\in\mathcal{A}\mid\mathcal{T}(\Phi)=\Phi\}.$ (14)
For $\Phi\in\mathcal{A}$, its restriction to $B\in\mathcal{P}_{\mathrm{fin}}(\mathbb{Z})$ is defined as
$\Phi|_{B}\coloneqq\sum_{A\subset B}\Phi(A)\in\mathcal{B}^{\otimes B},$ (15)
where we used the natural embedding
$\mathcal{B}^{\otimes A}\ni X\mapsto X\otimes I_{B\setminus A}\in\mathcal{B}^{\otimes B}$ (16)
(with $I_{C}$ the identity on $C\in\mathcal{P}_{\mathrm{fin}}(\mathbb{Z})$) to regard $\mathcal{B}^{\otimes A}\subset\mathcal{B}^{\otimes B}$. Accordingly, when viewing $\mathcal{B}^{\otimes B}=\oplus_{A\subset B}\mathcal{B}_{0}^{A}$, the projection onto each component is written as
$\mathcal{B}^{\otimes B}\ni Y\mapsto Y_{A}\in\mathcal{B}_{0}^{A}.$ (17)
III.2 Adjoint map and conserved quantities
[…]
$\operatorname{ad}_{\Phi}(\Psi)(A)\coloneqq[\Phi,\Psi](A)\coloneqq[\Phi|_{B},\Psi|_{B}]_{A},$ (18)
where $B\in\mathcal{P}_{\mathrm{fin}}(\mathbb{Z})$ is any sufficiently large finite set containing $A$. Since for $C,C^{\prime}\in\mathcal{P}_{\mathrm{fin}}(\mathbb{Z})$,
$[\Phi(C),\Psi(C^{\prime})]_{A}\neq 0\$$\Rightarrow\ C\triangle C^{\prime}\subsetneq A\subset C\cup C^{\prime}$
$(\Rightarrow C\cap A\neq\emptyset),$ (19)
the definition is independent of the choice of $B$. Here $C\triangle C^{\prime}$ denotes the symmetric difference $C\cup C^{\prime}\setminus(C\cap C^{\prime})$.
The map $\operatorname{ad}_{\Phi}$ is $\mathbb{C}$-linear in $\Psi$, and $\Phi\mapsto\operatorname{ad}_{\Phi}$ is also linear. Moreover, it satisfies the following properties:
  • $r([\Phi,\Psi])\leq r(\Phi)+r(\Psi)-1$,
  • $\Phi,\Phi]=0$,
  • $\Phi,\Psi]^{\dagger}=-[\Phi^{\dagger},\Psi^{\dagger}]$.
[…]
$\operatorname{ad}_{\Phi_{1}}\circ\operatorname{ad}_{\Phi_{2}}-\operatorname{ad}_{\Phi_{2}}\circ\operatorname{ad}_{\Phi_{1}}=\operatorname{ad}_{[\Phi_{1},\Phi_{2}]}.$ (20)
This follows because the value at each $A\in\mathcal{P}_{\mathrm{fin}}(\mathbb{Z})$ can be evaluated within a sufficiently large finite region, where ordinary operator commutators satisfy the Jacobi identity.
We now define conserved quantities in this framework.
Definition 3.
[…]
III.3 Boost operator and main theorem
Next, we define the boost operator [26, 27, 28].
Definition 4.
[…]
$\mathcal{T}(\mathfrak{B}_{\Phi})-\mathfrak{B}_{\Phi}=\Phi.$ (21)
A key property of boost operators is the following [30, 31].
Proposition 5.
[…]
Proof.
Take any $A\in\mathcal{P}_{\mathrm{fin}}(\mathbb{Z})$. For a sufficiently large $B\supset A$, we have
$\tau_{A}(1)\operatorname{ad}_{\mathfrak{B}_{\Phi}}(\Psi)(A)\tau_{A}(1)^{\dagger}$
$=\tau_{A}(1)[\mathfrak{B}_{\Phi}|_{B},\Psi|_{B}]_{A}\tau_{A}(1)^{\dagger}$
$=\left(\tau_{B}(1)[\mathfrak{B}_{\Phi}|_{B},\Psi|_{B}]\tau_{B}(1)^{\dagger}\right)_{A+1}$
$=\bigl[\tau_{B}(1)\mathfrak{B}_{\Phi}|_{B}\tau_{B}(1)^{\dagger},\,\tau_{B}(1)\Psi|_{B}\tau_{B}(1)^{\dagger}\bigr]_{A+1}$
$=\bigl[\mathcal{T}(\mathfrak{B}_{\Phi})|_{B+1},\,\Psi|_{B+1}\bigr]_{A+1}$
$=[\mathfrak{B}_{\Phi}|_{B+1},\Psi|_{B+1}]_{A+1}+[\Phi_{B+1},\Psi|_{B+1}]_{A+1}$
$=[\mathfrak{B}_{\Phi}|_{B+1},\Psi|_{B+1}]_{A+1}=\operatorname{ad}_{\mathfrak{B}_{\Phi}}(\Psi)(A+1),$ (22)
where in the last line we used that $\Psi$ is a conserved quantity of $\Phi$. ∎
Under these definitions, Theorem 1 can be restated as the following general theorem.
Theorem 6.
[…]
$\Phi^{(n)}\coloneqq(\operatorname{ad}_{\mathfrak{B}_{\Phi}})^{n-1}(\Phi)$ (23)
is a translationally invariant conserved quantity of $\Phi$. Furthermore, the family $\{\Phi^{(n)}\}_{n=1}^{\infty}$ consists of mutually commuting operators.
IV Proof of Main Theorem
For notational simplicity, we use the same notation as in Theorem 1 throughout the proof. The proof of Theorem 6 proceeds in exactly the same way as that of Theorem 1, upon replacing $\hat{H}$, $\hat{Q}^{(n+1)}$, and $\hat{B}_{\hat{H}}$ with $\Phi$, $\Phi^{(n)}$, and $\mathfrak{B}_{\Phi}$, respectively.
@@ IV Proof of Main Theorem
If $\hat{H},\hat{Q}^{(m+1)}]=0$ holds for all $m\geq 1$, then $\hat{Q}^{(k+1)},\hat{Q}^{(l+1)}]=0$ holds for $k,l\geq 1$.
@@
We proceed by induction on $M=k+l(\geq 2)$. For $M=2$, it is trivial that $\hat{H},\hat{H}]=0$ holds. Assume that the claim is valid up to $M=M^{\prime}-1\geq 2$, and consider $M=M^{\prime}$. From the Jacobi identity and the induction hypothesis, for $1\leq k\leq M^{\prime}-1$ with $l=M^{\prime}-k$, we have
$[\hat{Q}^{(k+1)},\hat{Q}^{(l+1)}]$ $=-[\hat{Q}^{(k)},\hat{Q}^{(l+2)}]+[\hat{B}_{\hat{H}},[\hat{Q}^{(k)},\hat{Q}^{(l+1)}]]$
$=-[\hat{Q}^{(k)},\hat{Q}^{(l+2)}]$
$=[\hat{Q}^{(k-1)},\hat{Q}^{(l+3)}]-[\hat{B}_{\hat{H}},[\hat{Q}^{(k-1)},\hat{Q}^{(l+2)}]]$
$=[\hat{Q}^{(k-1)},\hat{Q}^{(l+3)}]$
@@
$=(-1)^{k-1}[\hat{Q}^{(2)},\hat{Q}^{(M^{\prime})}]=0,$ (24)
where we used the assumption $\hat{H},\hat{Q}^{(M^{\prime})}]=0$ in the last equality. Thus the claim follows. ∎
From this lemma, it suffices to prove Theorem 1 by showing: If $\hat{H},\hat{Q}^{(3)}]=0$ holds, then $\hat{H},\hat{Q}^{(n)}]=0$ also holds for all $n\geq 4$. The key step is the following observation.
@@
Consider a finite system of size $N$ with the periodic boundary condition. Note that the commutativity of translationally invariant local operators is a local property, and hence independent of $N$. Thus, for sufficiently large $N$ compared with $n$, the assumption can be replaced with the finite-size relation $\hat{H}_{[N]},[\hat{H}_{[N]},\hat{A}^{(n)}_{[N]}]]=0$ (See Corollary 19 in Appendix A for a formal proof). By considering the Hilbert-Schmidt norm of $\hat{H}_{[N]},\hat{A}^{(n)}_{[N]}$], we obtain
@@
It should be noted that we cannot dispense with the assumption of translational invariance in Lemma 9, since $\hat{H},[\hat{H},\hat{B}_{\hat{H}}]]=0$ does not necessarily imply that $\hat{Q}^{(3)}=-[\hat{H},\hat{B}_{\hat{H}}]=0$.
We now summarize the proof of Theorem 1.
Proof of Theorem 1.
We prove by induction to show that for all $n\geq 2$, $\hat{Q}^{(n)}$ is a translationally invariant conserved quantity of $\hat{H}$. For $n=2$, the statement is trivial since $\hat{Q}^{(2)}=\hat{H}$.
Assume the claim holds for $n=m-1\geq 2$, and consider the case $n=m$. As stated in Eq. (4) (and Proposition 5), if $\hat{Q}^{(m-1)}$ is a translationally invariant conserved quantity, then $\hat{Q}^{(m)}$ is also translationally invariant. From Lemma 8, we have $\hat{H},[\hat{H},\hat{Q}^{(m)}]]=0$, and by Lemma 9, it follows that $\hat{H},\hat{Q}^{(m)}]=0$. This completes the proof. ∎
V Dichotomy on integrability